If X - Y = 9
If X + Z = 12
If Z - W = 14
If Y + W = 2
what's W X Y Z Values.

Answer by Alan3354(69443)

If X - Y = 9 --> x = y + 9
If X + Z = 12
If Z - W = 14
If Y + W = 2
----------
X + Z = 12 --> y+9 + z = 12 --> y + z = 3
Z - W = 14
Y + W = 2
----------
y + z = 3
Z - W = 14
Y + W = 2 --> w = -y + 2
---------
y + z = 3
Z - W = 14
w = -y + 2
---------
y + z = 3
Z - W = 14 --> z - (-y+2) = 14 --> y + z = 12
w = -y + 2
============
y + z = 3 and y + z = 12 --> inconsistent, no solution.